OpenStudy (smarty):

The second step to solving the equation 2x – 7 – (x – 5) = 0 is to

OpenStudy (smarty):

Is it combine like terms or add the opposite to both sides?

OpenStudy (anonymous):

Adding the opposite to both sides would only add a step later to get X on one side. You technically can't combine like terms because of the items in parentheses. You would get the items out of the parentheses like so: 2x-7 + (-1)(x-5) = 0 2x-7 + -x + 5 = 0 x-2=0 x=2

OpenStudy (smarty):

I was talking about the step after distributing @Wired